# Daniele Bartoli1, Mohit Pal2, Pantelimon Stănică3, Tommaso Toccotelli1

### v1

@@ Abstract
In this note we show (for a large enough dimension of the underlying field) a conjecture of [C. Beierle, C. Carlet, G. Leander, L. Perrin, A further study of quadratic APN permutations in dimension nine, Finite Fields Appl. 81 (2022), 102049] on a trivariate permutation. This function is a global representation of two new sporadic quadratic APN permutations in dimension $9$ found by [C. Beierle, G. Leander, New instances of quadratic APN functions, IEEE Trans. Inf. Theory 68(1) (2022), 670–678].
@@
Keywords: Finite fields, Permutation polynomials, Differential uniformity, varieties, irreducible components
Mathematics Subject Classification 2020: 12E20, 11T06
1 Introduction and tools from algebraic geometry
Let $q=2^{m},$ $m\in\mathbb{N},$ and denote by $\mathbb{F}_{q}$ the finite field with $q$ elements. For any positive integer $n$, we denote by $\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$, the ring of polynomials in $n$ indeterminates over finite field $\mathbb{F}_{q}$. An element $f\in\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$ is called a permutation polynomial in $n$ variables if the equation $f(X_{1},\ldots X_{n})=a$ has $q^{n-1}$ solutions in $\mathbb{F}_{q}^{n}$ for each $a\in\mathbb{F}_{q}$. Let $F:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ be a map given by
$F(X_{1},\ldots,X_{n})=(f_{1}(X_{1},\ldots,X_{n}),\ldots,f_{n}(X_{1},\ldots,X_{n})),$
where $f_{i}\in\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$, then $F$ is called a vectorial permutation if it induces a permutation on $\mathbb{F}_{q}^{n}$.
Vectorial Boolean functions are fundamental building blocks in symmetric cryptography, since many block ciphers employ these as components in their S-boxes. The security of a block cipher depends upon the immunity of its substitution boxes against various kinds of cryptographic attacks. For instance, a low differential uniformity [13] to resist the differential attacks [5]. For a prime $p$ and positive integer $n>0$, the differential uniformity of an $(n,n)$-function $F:\mathbb{F}_{p^{n}}\to\mathbb{F}_{p^{n}}$ is defined as the maximum number of solutions $x$ of the differential equation $F(x+a)-F(x)=b$, where $a\neq 0,b\in\mathbb{F}_{p^{n}}$. When then differential uniformity is $1$, we say that the function is perfect nonlinear (or planar). It may be noted that planar functions exist only in the case of odd characteristic. The lowest possible differential uniformity of functions over finite fields of even characteristic is $2$ and such functions are called almost perfect nonlinear (APN). APN functions in characteristic $2$ have been studied since the 90’s and a lot of research has been done in this direction in recent years (see [7] and references therein).
For the necessary algebraic geometry background, we refer to [9, 10]. We use the notations $\mathbb{P}^{r}(\mathbb{F}_{q})$ and $\mathbb{A}^{r}(\mathbb{F}_{q})$ (or $\mathbb{F}_{q}^{r}$) to denote the projective and the affine space of dimension $r\in\mathbb{N}$ over the finite field $\mathbb{F}_{q}$, respectively. A variety, and more specifically a curve or a surface (i.e. a variety of dimension 1 or 2, respectively), is described by a certain set of equations with coefficients in $\mathbb{F}_{q}.$ We say that a variety $\mathcal{V}$ is absolutely irreducible if there are no varieties $\mathcal{V}^{\prime}$ and $\mathcal{V}^{\prime\prime}$ defined over the algebraic closure of $\mathbb{F}_{q}$ (denoted by $\overline{\mathbb{F}_{q}}$) and different from $\mathcal{V}$ such that $\mathcal{V}=\mathcal{V}^{\prime}\cup\mathcal{V}^{\prime\prime}$. If a variety $\mathcal{V}\subseteq\mathbb{A}^{r}(\mathbb{F}_{q})$ is defined by $F_{i}(X_{1},\dots,X_{r})=0,$ for $i=1,\dots s,$ an $\mathbb{F}_{q}$-rational point of $\mathcal{V}$ is a point $(x_{1},\dots,x_{r})\in\mathbb{A}^{r}(\mathbb{F}_{q})$ such that $F_{i}(x_{1},\dots,x_{r})=0,$ for $i=1,\dots,s.$ The set of the $\mathbb{F}_{q}$-rational points of $\mathcal{V}$ is usually denoted by $\mathcal{V}(\mathbb{F}_{q}).$ If $s=1,$ $\mathcal{V}$ is called a hypersurface and it is absolutely irreducible if the corresponding polynomial $F(X_{1},\dots,X_{r})$ is absolutely irreducible, i.e. it possesses no non-constant factors over $\overline{\mathbb{F}_{q}}.$ Moreover, we say that $\mathcal{V}$ is a variety of degree $d$ (and write $\deg(\mathcal{V})=d$) if $d=\#(\mathcal{V}\cap H),$ where $H\subseteq\mathbb{A}^{r}(\overline{\mathbb{F}_{q}})$ is a general projective subspace of dimension $r-s.$ To determine the degree of a variety is generally not straightforward; however an upper bound to $\deg(\mathcal{V)}$ is given by $\prod_{i=1}^{s}\deg(F_{i}).$ We also recall that the Frobenius map $\Phi_{q}:x\mapsto x^{q}$ is an automorphism of $\mathbb{F}_{q^{k}}$ and generates the group $Gal(\mathbb{F}_{q^{k}}/\mathbb{F}_{q})$ of automorphisms of $\mathbb{F}_{q^{k}}$ that fixes $\mathbb{F}_{q}$ pointwise. The Frobenius automorphism induces also a collineation of $\mathbb{A}^{r}(\overline{\mathbb{F}_{q}})$ and an automorphism of $\overline{\mathbb{F}_{q}}[X_{1},\dots,X_{r}].$
A crucial point in our investigation of permutation trinomials over $\mathbb{F}_{q^{3}}$ is to prove the existence of suitable $\mathbb{F}_{q}$-rational points in algebraic surfaces $\mathcal{V}$ attached to each permutation trinomial. This is reached by proving the existence of absolutely irreducible $\mathbb{F}_{q}$-rational components in $\mathcal{V}$ and lower bounding the number of their $\mathbb{F}_{q}$-rational points. To this end, generalizations of Hasse-Weil type bounds for algebraic varieties are needed. To ensure the existence of a suitable $\mathbb{F}_{q}$-rational point of $\mathcal{V},$ we need the following result.
Theorem 1.
@@ 1 Introduction and tools from algebraic geometry
[8, Theorem 7.1] Let $\mathcal{V}\subseteq\mathbb{A}^{n}(\mathbb{F}_{q})$ be an absolutely irreducible variety defined over $\mathbb{F}_{q}$ of dimension $r>0$ and degree $\delta.$ If $q>2(r+1)\delta^{2},$ then the following estimate holds:
|#$\big{|}\#(\mathcal{V}(\mathbb{A}^{n}(\mathbb{F}_{q})))-q^{r}\big{|}\leq(\delta-1)(\delta-2)q^{r-1/2}+5\delta^{13/3}q^{r-1}.$ (1)
In our approach we will make use of the following result.
Lemma 2.

### v2

# Non-permutation phenomena in trivariate families over $\mathbb{F}_{2^{m}}$ and resolution of a conjecture

@@ Abstract
Constructing permutation polynomials over finite fields, particularly those with simple algebraic structure in multiple variables, is a fundamental problem with applications in cryptography and coding theory. Recently, Li and Kaleyski (IEEE Trans. Inf. Theory, 2024) generalized two sporadic quadratic APN permutations into infinite families of trivariate functions. Motivated by their work, we investigate conditions under which generalized trivariate functions fail to be permutations. We establish necessary conditions on coefficient parameters that prevent the permutation property, provide a complete computational classification for small field extensions, and prove general non-permutation results. As a key application of our algebraic geometry approach, we resolve the permutation part of a conjecture by Beierle, Carlet, Leander, and Perrin (Finite Fields Appl., 2022) regarding a related trivariate form. Specifically, we prove that for all odd characteristic-2 extension degrees $m\geq 23$, their function $C_{u}$ is not a permutation over $\mathbb{F}_{2^{m}}^{3}$ for any $u\in\mathbb{F}_{2^{m}}^{*}$, resolving the permutation part of their conjecture for sufficiently large fields.
@@
Keywords: finite fields, permutation polynomials, vectorial Boolean functions, algebraic geometry, algebraic varieties
Mathematics Subject Classification 2020: 12E20, 11T06, 14G15
1 Introduction
Let $q=2^{m}$ with $m\geq 1$, and let $F:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{n}$ be a vectorial function. When $F$ is bijective, it is a permutation and is a natural S-box candidate in symmetric cryptography. For even characteristic, differential behavior is encoded by the equations
$F(x+a)+F(x)=b,\qquad a\in\mathbb{F}_{q}^{n}\setminus\{0\},\ b\in\mathbb{F}_{q}^{n}.$
Low differential uniformity is essential against differential cryptanalysis [5]. In particular, APN functions (uniformity $2$) are optimal in characteristic two.
Two sporadic APN permutations over $\mathbb{F}_{2^{9}}$ discovered by Beierle and Leander [4] triggered several trivariate generalizations. Beierle, Carlet, Leander, and Perrin [3] investigated
$C_{u}(X,Y,Z)=\big(X^{3}+uY^{2}Z,\ Y^{3}+uXZ^{2},\ Z^{3}+uX^{2}Y\big),$
and conjectured that, beyond the known small-dimensional cases, this form should not yield permutations/APN functions. The APN part was settled in [2], while the permutation part remained open. Li and Kaleyski then introduced two infinite trivariate APN families with a different quadratic organization [10].
This paper is a structural non-permutation study. Our objective is not only to settle one conjecture case, but to isolate mechanisms that force non-permutation behavior in broad trivariate classes.
Main contributions.
  1. 1.
    We start from an $8$-parameter trivariate family and prove that every non-degenerate instance is affine-equivalent to a canonical $4$-parameter model (Proposition 3.1).
  2. 2.
    For the canonical model, we derive explicit algebraic conditions whose solvability implies non-permutation (Theorems 4.1, 4.2, 4.3).
  3. 3.
    For $m=3$, we obtain a complete characterization by exhaustive computation: exactly 2020 parameter choices are permutations, split into two structural families (Proposition 5.1, Table 1).
  4. 4.
    As an application of the same geometry framework, we resolve the permutation part of the Beierle–Carlet–Leander–Perrin conjecture for large odd dimensions: for every odd $m\geq 23$ and every $u\in\mathbb{F}_{2^{m}}^{*}$, $C_{u}$ is not a permutation (Theorem 6.3).
The remainder of the paper is organized as follows. Section 2 collects geometric ingredients (Lang–Weil type bounds and component arguments). Section 3 introduces the full trivariate family and proves affine reduction to the canonical form. Section 4 develops the differential-system method and derives the necessary non-permutation criteria. Section 5 gives the complete $m=3$ classification. Section 6 applies the framework to the Beierle conjecture instance. Section 7 closes with consequences and open directions.
2 Preliminaries and algebraic geometry tools
We collect notation and algebraic-geometric tools used in the sequel in one continuous setup. The derivative notation is fixed by $D_{a}F(x)=F(x+a)+F(x)$ for $a\neq 0$.
We also invoke the standard finite-field point-count and elimination facts in the exact forms needed later, so each reduction step is explicit inside the proofs. Unless explicitly stated otherwise, we work over $\mathbb{F}_{2^{m}}$ with $m$ odd and $m\geq 3$, and set $q=2^{i}$ with $\gcd(i,m)=1$. All parameters are elements of $\mathbb{F}_{2^{m}}$. Whenever nondegeneracy conditions are required (for instance, nonvanishing coefficients in normalized forms), they are stated explicitly in each theorem.
For a polynomial system $\mathcal{S}$, by its solution set we mean the affine set in the ambient space over the specified ground field.
We employ standard tools from algebraic geometry; comprehensive background appears in [8, 9]. We denote by $\mathbb{P}^{r}(\mathbb{F}_{q})$ and $\mathbb{A}^{r}(\mathbb{F}_{q})$ the projective and affine spaces of dimension $r$ over $\mathbb{F}_{q}$, with $\overline{\mathbb{F}_{q}}$ denoting the algebraic closure.
A variety $\mathcal{V}$ is the common zero set of finitely many polynomials. An affine $\mathbb{F}_{q}$-rational variety $\mathcal{V}\subset\mathbb{A}^{r}(\overline{\mathbb{F}_{q}})$ is defined by polynomials $F_{1},\ldots,F_{s}\in\mathbb{F}_{q}[X_{1},\ldots,X_{r}]$. The set $\mathcal{V}(\mathbb{F}_{q})=\mathcal{V}\cap\mathbb{A}^{r}(\mathbb{F}_{q})$ consists of $\mathbb{F}_{q}$-rational points. A variety is absolutely irreducible if it cannot be written as a union of proper subvarieties over $\overline{\mathbb{F}_{q}}$.
The dimension of a variety is the maximal length of a chain of distinct, nonempty, absolutely irreducible subvarieties. Varieties of dimension 1, 2, and r1r-1 are called curves, surfaces, and hypersurfaces. The degree of an $s$-dimensional projective variety is the number of intersection points with a general projective subspace of complementary dimension.
The Frobenius map $\Phi_{q}:x\mapsto x^{q}$ fixes $\mathbb{F}_{q}$ pointwise and induces automorphisms on polynomial rings and projective/affine spaces. A key tool is finding $\mathbb{F}_{q}$-rational points on algebraic surfaces. This requires identifying absolutely irreducible $\mathbb{F}_{q}$-rational components and applying point-counting estimates:
Theorem 2.1 (Strengthened Lang-Weil [7, Theorem 7.1]).
@@ 2 Preliminaries and algebraic geometry tools
Let $\mathcal{V}\subseteq\mathbb{A}^{n}(\overline{\mathbb{F}_{q}})$ be an absolutely irreducible $\mathbb{F}_{q}$-rational variety of dimension $r>0$ and degree $\delta$. If $q>2(r+1)\delta^{2}$, then
$\big|\#\mathcal{V}(\mathbb{F}_{q})-q^{r}\big|\leq(\delta-1)(\delta-2)q^{r-1/2}+5\delta^{13/3}q^{r-1}.$
We will also make use of the following result.
Lemma 2.2 ([1, Lemma 2.1]).